OpenStudy (anonymous):

solve for h, show your work 4 (h + 2) = 3 (h - 2)

OpenStudy (anonymous):

4h + 8 = 3h - 6 what do you do after distribution?

OpenStudy (anonymous):

Take all the "h" terms to one side and the numbers to the other. Keep in mind that everything changes signs when you switch side,s what do you get?

OpenStudy (anonymous):

4h - (-3h) = 6 - (-8) ?

OpenStudy (anonymous):

Eh...no. 4(h + 2) = 3(h -2) 4h + 8 = 3h - 6 4h - 3h = -6 -8 h = ?

OpenStudy (anonymous):

h = -14

OpenStudy (anonymous):

Very good.

OpenStudy (anonymous):

my math teacher doesnt explain things to me in class ): when 3h switches sides, why doesnt it turn into a negative?

OpenStudy (anonymous):

It does. 4(h + 2) = 3(h -2) 4h + 8 = 3h - 6 4h - 3h = -6 -8 h = ? Look that 3h became -3h when we brought it to the left side.

OpenStudy (anonymous):

okay. youve helped me alot tonight, i cant thank you enough